OpenStudy (anonymous):

how many real-number solutions does the equation have? –7x² + 6x + 3 = 0

OpenStudy (bahrom7893):

Look at the discriminant

OpenStudy (bahrom7893):

D=b^2-4ac=36+28*3=36+84=120>0, therefore it has two real solutions